# Do we really have to filter out random noise in pre-training data for language models?

### v1

@@ 3.2 Results
These positive experimental outcomes further corroborate the robustness of language models and provide insights into why pre-training on large-scale datasets that inevitably contain significant amounts of noise can still yield high-performing models. These somewhat unexpected findings naturally prompt us to explore the underlying reasons.
@@ 3.3 Theoretical Analysis
In the analysis below, we focus on the impact of random noise on NTP loss, as pre-training loss is crucial for the performance on downstream tasks (Saunshi et al., 2021; Wei et al., 2021; Liu et al., 2023; Zheng et al., 2023a). Specifically, we are interested in the difference of NTP Loss between a model $h^{*}$ trained on a noise-free dataset and a model $h$ trained with a noisy dataset. We begin by noting that sampling from the clean distribution should not yield random gibberish and vice versa. Mathematically, this implies that for any prefix $r$ sampled from $P^{n}_{X}$, the probability under the clean distribution $P^{c}_{X}(r)$ is zero. Thus, we make the following assumption:
Assumption 1.
$P^{c}$ and $P^{n}$ have disjoint support sets, i.e., ${\text{supp}}(P^{c})\cap{\text{supp}}(P^{n})=\emptyset$.
The subsequent proposition demonstrates that the error $\epsilon$ introduced to the loss due to random noise is less than the proportion $\alpha$ of random noise in the dataset.
@@
With these theoretical results in hand, we then conduct multiple experiments to substantiate their validity. First, we plot the trend of NTP loss on random noise within the training set throughout the learning process, as shown in Figure 3(a). It is evident that the loss on random noise decreases at a very slow rate, indicating that the model struggles to efficiently learn the distribution of random noise.
Next, we add 5% Gaussian-distributed noise to the training dataset and compare the results with models trained on 5% random noise. Specifically, we replace the uniform distribution mentioned above with a Gaussian distribution characterized by a mean of 25128 and standard deviation of 500. As depicted in Figure 3(b), the loss on Gaussian noise is lower than that on the random noise. According to Proposition 1, since the Gaussian distribution corresponds to a high $p_{n}$, we can predict that a model trained on Gaussian noise will exhibit a lower loss on $P^{c}$. Figure 3(c) confirms our prediction, thus further validating the proportions.
@@
Table 1: Accuracy on 4 text classification benchmark. 0% represents a model trained on $P^{c}$, 1% and so on denote the proportion of random noise, and Gaussian refers to Gaussian noise. cited from Saunshi et al. (2021).
@@
3.4 Experiments on Other Text Corpus
To further investigate the impact of random noise on model generalization, we evaluate the next-token prediction loss of the trained models on data crawled from arXiv and Wikipedia. The results are illustrated in Figure 4. Surprisingly, models trained with added noise outperformed those trained on $P^{c}$. This counterintuitive finding aligns with previous work in visual domains (Zada et al., 2022), suggesting that incorporating random noise into training sets might enhance model robustness. Additionally, we observe that the performance of models subjected to Gaussian noise varies across different datasets. These observations warrant further investigation.
3.5 Broader Impact of the Results
One immediate direction is the training of multilingual models (Pires et al., 2019; Chi et al., 2020; Yang et al., 2024a). Clearly, tokens corresponding to different languages are distinct, and their distributions naturally satisfy Assumption 1. For example, in an English-French bilingual model, let $P^{c}$ represent English and $P^{n}$ represent French. Supposing the pre-training corpus consists of an equal distribution of English and French, and given that the two distributions are similar, we can assume that $p_{c}\approx p_{n}$, leading to $\epsilon\approx 0$. This provides a theoretical foundation for the success of multilingual models. See Appendix D.3 for more details.
Beyond language modality, random white noise has received increased attention in the speech domain (Chen et al., 2022, 2021; Yang et al., 2024c, 2023b). Since our theory applies to any cross-entropy-like loss, it can also explain why speech models pre-trained on very noisy large-scale datasets, such as Gigaspeech (Chen et al., 2021), which contain significant background noise and prolonged silence at the beginning and end of a few samples, still perform remarkably well.
4 Reducing the Noise with Local Gradient Matching
In Section 3, we know that the influence of noise on NTP loss is rather small. However, Figure 3(c) and Table 1 show that the Gaussian noise-trained model with lower NTP loss suffers a 1.5% decrease in accuracy in downstream tasks. Although this might be mitigated during the pre-training phase, considering that most practitioners fine-tune pre-trained models rather than training them from scratch, we propose a novel black-box fine-tuning method termed Local Gradient Matching loss to tame the influence. Extensive experiments across 8 natural language understanding and 14 image classification benchmark datasets further demonstrate that the proposed method consistently enhances performance across different backbones and modalities. We also provide a theoretical analysis.
@@ 4 Reducing the Noise with Local Gradient Matching
4.1 Method
In the preceding analysis, we demonstrate that the population-level loss function is only marginally affected by random noise. However, during the SGD training process, its presence introduces certain noise into the gradients. Prior studies (Chen et al., 2023; Xie et al., 2021c) have shown that artificially added gradient noise can hurt the model’s generalization. Therefore, inspired by Sharpness-Aware Minimization (SAM) (Foret et al., 2021; Zhang et al., 2023; Zhao et al., 2022; Wen et al., 2023) and noise-robust fine-tuning methods (Hua et al., 2023, 2021; Jiang et al., 2020), we propose explicitly enhancing the denoising capabilities of the downstream task head by aligning local gradients.
@@ 4.1 Method
(a) No regularization.
(b) With $L^{2}$ regularization.
(c) With $\mathcal{L}_{gm}$.
Specifically, let $C$ denote the number of classes in the downstream task, and let $g_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ represent the linear or MLP classification head parameterized by $\theta$. Let $t^{*}$ be the feature extracted by $h^{*}$, $t$ be the feature extracted by $h$, and $y$ be the corresponding label. Let $\ell(\hat{y},y)$ be the loss function(typically cross-entropy), and $\mathcal{L}_{ce}(\mathcal{D},g_{\theta})=\mathbb{E}_{(t,y)\sim\mathcal{D}}\ell(g_{\theta}(t),y)$ be the population-level loss where $\mathcal{D}$ represents the joint distribution of downstream features and labels. Due to the additional randomness introduced by $h$ as a result of noise, $t$ can be viewed as $t^{*}$ perturbed by minor disturbances. If both $t^{*}$ and $t$ were known, their distribution could be aligned to achieve denoising. However, in practical applications, it is challenging to obtain $t^{*}$. To construct contrastive sample pairs without $t^{*}$, we add Gaussian noise to $t$ to obtain $\hat{t}$:
$\hat{t}=t+\gamma\cdot\delta\vspace{-2mm}$ (3)
$\mathcal{L}_{gm}(\theta)$ $=||\mathbb{E}_{(t,y)\sim\mathcal{D}}\nabla_{\theta}\ell(g_{\theta}(t),y)$ (4)
$\quad-\mathbb{E}_{(\hat{t},y)\sim\hat{\mathcal{D}}}\nabla_{\theta}\ell(g_{\theta}(\hat{t}),y)||_{2}$
@@
Table 4: Evaluation of our method combined with SOTA fine-tuning techniques utilizing BERT-Large as the backbone model across 4 datasets. cited from Hua et al. (2021)
4.2 Theoretical Analysis
Proposition 2.
Proposition 2 demonstrates that $\mathcal{L}_{gm}$ is closely associated with the smoothness of the loss function in both the parameter space and the input space. As a flat minima is widely acknowledged to benefit the generalization of neural networks (Xie et al., 2021b; Baldassi et al., 2021), it explains the effectiveness of $\mathcal{L}_{gm}$. The final loss function is:
$\mathcal{L}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{gm}\vspace{-2mm}$ (6)
4.3 Experiments
We first conduct extensive experiments using trained GPT-2 models. Then, to further validate the novelty and effectiveness of the LGM loss, we conduct additional experiments using Llama-3 (Dubey et al., 2024) and vision models. These experiments are intended to showcase the generalizability of our approach beyond the specific context of GPT-2, demonstrating its applicability across different types of models and tasks. While these experiments enrich our study, they are not the core focus but rather supplementary evidence supporting the broader applicability of our proposed solution. Details can be found in Appendix E.
We validate the performance of $\mathcal{L}_{gm}$ on models pre-trained with noisy data using four commonly used classification datasets: SST-2, SST-fine, 20newsgroup, and CR. The training hyperparameters follow those of Saunshi et al. (2021), where $\gamma=0.01$ and $\lambda=0.15$ apply to all four experiments. In line with the approach described by Chen et al. (2024), we freeze the model parameters and only fine-tune a linear or MLP classifier head. As shown in Table 1, our model achieves competitive results without reaching the number of training iterations of GPT-2, and $\mathcal{L}_{gm}$ consistently boosts performance.
@@ 4.3 Experiments
Table 5: Ablation Study. To investigate the effects of reducing $\mathcal{L}_{gm}$, experiments are conducted to examine the impact of separately reducing the norm versus increasing the cosine similarity.
@@
0.001 0.001 76.31 78.54
0.05 0.05 76.54 79.51
0.1 0.1 76.43 79.12
@@ 5 Conclusion
In this paper, we investigate the random noise present in language model pre-training datasets, which is inevitable in real-world scenarios but receives little attention. We pre-train multiple GPT-2 models under varying noise levels and find that random noise has a minor impact on the pre-training loss. We then provide a theoretical explanation for this phenomenon and discover that our theory can elucidate the success of multilingual models. Interestingly, we observe that slight noise can sometimes enhance a model’s generalization ability. Then, building on the noisy model learning setup, we propose a novel local gradient matching loss. Extensive experiments across multiple datasets in both language and vision tasks, as well as with various backbone models, validate the effectiveness of our proposed method. We hope this work inspires more researchers to focus on data-centric AI.
Limitations
In this section, we discuss the limitations of this paper.
Firstly, due to limitations in computational resources and costs, we pre-train only the GPT-2 124M and 774M model(see Appendix D.5) on the OpenWebText dataset. Compared to today’s large language models, both the scale of OpenWebText and that of GPT-2 are relatively small. Additionally, the types of noise considered are limited to uniform and Gaussian distributions. However, based on Proposition 1, we argue that training GPT-2 on the Synthetic OpenWebText dataset is sufficient to uncover the essence of the issue, as Proposition 1 makes no assumptions about data distribution or model architecture.
@@ B.1 Explanation of Equation (1)
First, define a new probability measure $P_{3}$ as $P_{3}(A)=\alpha P_{1}(A)+(1-\alpha)P_{2}(A)$ for any measurable set $A\subseteq\mathcal{M}$. Here, $P_{3}$ is a convex combination of $P_{1}$ and $P_{2}$, and thus $P_{3}$ is also a valid probability measure (Le Gall, 2022).
@@ B.3 Justification of Proposition 1
(1) Numerous studies on compressing large language models, such as pruning (Wang et al., 2020; Kurtic et al., 2022; Zhang et al., 2024), quantization (Zhao et al., 2024; Jin et al., 2024; Liu et al., 2024), and distillation (Dasgupta et al., 2023; Hinton, 2015), have demonstrated that there exists a significant amount of redundancy within the parameters of large models. Therefore, we could first train a model on $P^{c}$ and then compress it, fine-tuning the surplus parameters on $P^{n}$. This approach would allow us to improve $p_{n}$ without altering $p_{c}$.
(2) A small proportion of data corresponding to $P^{n}$ can be introduced into $P^{c}$, making sure that $\alpha$ is extremely small. According to domain adaptation theory (Ben-David et al., 2010), this would only slightly increase $\mathcal{L}_{ntp}(P^{c})$. However, existing results (Shliazhko et al., 2024; Pires et al., 2019; Chi et al., 2020) indicate that pre-trained models like BERT or GPT on English text can exhibit strong multilingual capabilities with just a very limited amount of data. Consequently, compared to a model trained solely on $P^{c}$, the resulting model has a minor difference in $p_{c}$ but a relatively higher $p_{n}$.
@@
Definition 1 ($\beta$-smooth (Zheng et al., 2023b)).
@@
where $B(0,\rho)=\{\delta^{\prime}:||\delta^{\prime}||_{2}<\rho\}$ is a open ball.
@@ Appendix C Detailed Related Works
Data selection for language model training. LLMs (Yang et al., 2023c; Zhuang et al., 2024c, d; Xie et al., 2024; Pan et al., 2024; Weng et al., 2025; Fan et al., 2024c) have fundamentally change the landsape of current AI research (Pan et al., 2023; Fan et al., 2024a; Zhuang et al., 2024a, e; Yang et al., 2023a, 2024b). Large text corpora form the backbone of language models, with data quality being fundamental to their success and safety (Yang et al., 2023e, f). Elazar et al. (2024) conducted a systematic analysis of open-source text datasets such as The Pile (Gao et al., 2020) (used to train Pythia), C4 (Raffel et al., 2020) (used to train T5) and RedPajama (used to train LLaMA), revealing that they contain a significant amount of duplicate, toxic, synthetic, and low-quality content. Therefore, it is of great importance to thoroughly understand the impact of low-quality data within these pre-training datasets on the model’s performance, reliability, and safety. Allen-Zhu and Li (2024a, b) systematically investigated the effect of low-quality data and found that such data can significantly reduce the model’s knowledge capacity, sometimes by up to 20 times. Another research direction primarily focuses on the synthetic data of large language models, specifically examining the impacts of using data generated by LLMs for recursive training. The study by Shumailov et al. (2024) was the first to explore this issue and introduced the concept of "model collapse", indicating that recursive training can lead to the loss of information in tail tokens, ultimately resulting in the model producing nonsensical content. Seddik et al. (2024) mainly provided a theoretical explanation for why model collapse occurs, supporting their arguments with experimental evidence. Consequently, the importance of data selection cannot be overstated. Given that data selection is an NP-hard problem in terms of combinatorial optimization (Xiao et al., 2022), numerous heuristic algorithms have been proposed to expedite the process. Longpre et al. (2024) provided a comprehensive study on pre-training data selection and optimal ratios, offering practical recommendations. Yang et al. (2023d) proposed dataset pruning, an approach that assesses the impact of omitting training samples on model generalization and creates a minimal training subset with a controlled generalization gap. Chai et al. (2024) evaluated the impact of individual training samples on the dynamics of GPT model training. Li et al. (2024b) introduced the Instruction-Following Difficulty metric to assess the quality of instruction-tuning data. Xie et al. (2023b) employed importance resampling for data selection. Xie et al. (2023a); Lee et al. (2023) advocated for optimizing data composition and diversity. Despite these notable studies on data selection, they generally acknowledge that dataset noise degenerates model performance but lack a detailed understanding of how and to what extent, particularly in the case of random noise which is inevitable in large-scale datasets. Although Cherepanova and Zou (2024) investigated the influence of gibberish input, the random noise within the pre-training dataset is still underexplored. This paper aims to bridge the gap.
@@
Learning from Noisy Distributions. The majority of machine learning algorithms assume that training and test samples are independently and identically distributed (i.i.d.), a condition that is often not met in real-world scenarios. For instance, LLMs are pre-trained on datasets with all kinds of noise while their performance is evaluated by the user whose distribution is usually clean and meets real-world scenarios, which violates the i.i.d. assumption. Domain adaptation (Li et al., 2024a; Meng et al., 2022; Ma et al., 2023) addresses this issue when the distribution of the training data differs from that of the test data. Although domain adaptation methods attempt to reduce the statistical distribution discrepancy (Du and Li, 2023; Wu et al., 2024) or employ adversarial training (Li et al., 2021; Ru et al., 2024) to minimize the gap between source and target domains, they typically require access to unlabeled test data under a semi-supervised learning setup, which is impractical for LLM training. Another reason domain adaptation cannot be directly applied here is that domain adaptation theory (Ben-David et al., 2010) focuses on the performance of a model trained on one distribution when it is applied to another different but related distribution. This kind of bounds can be easily derived by Lemma 1. However, what we aim to investigate here is the extent of performance loss when comparing a model trained on one distribution (noisy dataset) to a model trained on another distribution (clean dataset).
@@
Apart from domain adaptation, there has been extensive research directly investigating noisy training sets. Noisy label learning Song et al. (2022); Lukasik et al. (2020) have explored the impact of incorrect labels on model performance. Regarding input feature noise, Smilkov et al. (2017) added perturbations to individual image inputs to enhance model interpretability, and Zada et al. (2022) added white noise image into the training dataset to tackle the class imbalance problem. However, most of these efforts have concentrated on image classification and do not consider the pre-training paradigm.
@@
Fine-tuning Pre-trained Models. The approach of initially pre-training model weights on large-scale datasets and subsequently fine-tuning them with downstream data has become the de facto standard in the fields of computer vision (Ye and Gao, 2024; Zhuang et al., 2025b; Fan et al., 2024b; Zhuang et al., 2025a), natural language processing (Wei et al., 2021; Xie et al., 2021a; Zhuang et al., 2024f; Tang et al., 2024) and audio processing (Yang et al., 2023b; Zhuang et al., 2024b, 2025c). For instance, Hua et al. (2023, 2021) proposed enhancing the performance of models by increasing their resistance to minor perturbations in intermediate layers. Meanwhile, Jiang et al. (2020) improved model robustness by adding regularization terms. Besides full-parameter fine-tuning, numerous parameter-efficient fine-tuning algorithms have been extensively studied. Zhang et al. (2021) introduced adapters into the original model architecture, optimizing only these parameters during fine-tuning. Zhou et al. (2022b, a) efficiently fine-tuned CLIP models (Radford et al., 2021) using learnable soft prompts. Hu et al. (2022) optimized models through learning low-rank residual weights. These methods achieved performance close to that of full-parameter fine-tuning while maintaining the generalization ability of the original models. However, they all require access to the model’s weights and loading them into GPU memory, which can be challenging for today’s large models, especially when state-of-the-art models’ parameters are not publicly available. Therefore, in this paper, we follow the NML setup and explore efficient ways to fine-tune the downstream task head under a black-box scenario.
@@
Implicit Regularization and Sharpness-aware Minimization. Achieving good generalization in neural networks optimized using gradient descent algorithms has long been a research focus in deep learning theory. Barrett and Dherin (2021) explored the properties of stochastic gradient descent (SGD), finding that SGD implicitly constrains the gradient norm. Based on this observation, Sharpness-aware minimization (SAM) (Zhang et al., 2023; Foret et al., 2021; Wen et al., 2023; Xie et al., 2023c) improves generalization by incorporating the gradient norm as a regularization term. Our method can be seen as drawing inspiration from SAM but differs in that our optimization objective is the model’s resilience to input noise rather than seeking flat minima in the parameter space.

### v2

@@ 3.3 Experiments
As illustrated in Figure 3(a), the impact of random noise remains negligible even when the model size scales up to 2.7B parameters. Furthermore, when evaluating the trained model on Wikipedia and arXiv datasets for NTP loss (detailed in Appendix D.3), we observe that random noise can even be beneficial in certain instances. These counter-intuitive experimental outcomes further corroborate the robustness of language models and provide insights into why pre-training on large-scale datasets that inevitably contain significant amounts of noise can still yield high-performing models. These somewhat unexpected findings naturally prompt us to explore the underlying reasons.
@@ 3.4 Theoretical Analysis
In the analysis below, we focus on the impact of random noise on NTP loss which is crucial for the performance on downstream tasks [24, 25, 26, 27]. Specifically, we are interested in the difference of NTP Loss between a model $h^{*}$ trained on a noise-free dataset and a model $h$ trained with a noisy dataset.
@@
With these theoretical results in hand, we then conduct multiple experiments to substantiate their validity. First, we plot the trend of NTP loss on random and Gaussian noise within the training set throughout the learning process, as shown in Figure 3(b). It is evident that the loss decreases at a very slow rate, indicating that the model struggles to efficiently learn the distribution of noise. Furthermore, the loss on Gaussian noise is lower than that on the random noise. According to Proposition 1, since the Gaussian distribution corresponds to a high $p_{n}$, we can predict that a model trained on Gaussian noise will exhibit a lower loss on $P^{c}$. Figure 3(c) confirms our prediction, thus validating the proportions.
3.5 Generalized Impact beyond Random Noise
One immediate direction is the training of multilingual models [28, 29, 30]. Apparently, tokens corresponding to different languages are usually distinct, and the sentences in one language usually do not appear in another language. Therefore, the distributions of different languages naturally satisfy Assumption 1. For example, in an English-Chinese bilingual model, let $P^{c}$ represent English and $P^{n}$ represent Chinese. Supposing the pretraining corpus consists of an equal distribution of English and Chinese, and given that the two distributions are natural languages, we can assume that $p_{c}\approx p_{n}$, leading to $\epsilon\approx 0$. This provides a theoretical foundation for the success of multilingual models. See Appendix D.4 for more details.
Beyond the language modality, we demonstrate that our theoretical framework naturally extends to emerging native multimodal autoregressive large models (MLLMs). For example, VARGPT [15, 31] uses a visual tokenizer for image understanding and generation, while UniAudio [16] achieves state-of-the-art audio generation by modeling the joint distribution of audio and text tokens. Taking audio foundation models as a representative case, they typically employ a neural codec [32] to discretize continuous audio signals into token sequences. The multimodal input is then formatted as a concatenated sequence with text tokens before audio tokens, enabling standard next-token prediction during pretraining to learn conditional audio generation from textual prompts.
This paradigm inherently requires simultaneous modeling of text and audio token distributions. Apparently, the sets of audio tokens and text tokens are disjoint, thereby satisfying Assumption 1. Therefore, Proposition 1 provides a theoretical justification for the empirical observation that joint distribution modeling does not compromise model performance. For example,UniAudio has reported that although they mask text tokens during NTP pretraining to focus solely on audio prediction, empirical results show no significant improvement in loss metrics compared to the full joint prediction approach. This alignment between theoretical predictions and empirical observations underscores the practical relevance of our framework for analyzing modern multimodal architectures.
4 Reducing the Noise with Local Gradient Matching
Table 1: Accuracy on 4 text classification benchmark. 0% represents a model trained on $P^{c}$, 1% and so on denote the proportion of random noise. cited from [24].
@@ 4 Reducing the Noise with Local Gradient Matching
In Section 3, we know that the influence of noise on NTP loss is rather small. However, Figure 3(c) and Table 1 show that the Gaussian noise-trained model with lower NTP loss suffers a 1.5% decrease in accuracy in downstream tasks. Therefore, in order to tame the potential influence in downstream tasks, we propose a novel black-box fine-tuning method termed Local Gradient Matching loss. Extensive experiments across 8 natural language understanding and 14 image classification benchmark datasets further demonstrate that the proposed method consistently enhances performance across different backbones and modalities. We also provide a theoretical analysis.
It is important to emphasize that the rationale for investigating pre-training in Section 3 while proposing a fine-tuning-based approach for downstream tasks in this section stems from current conventions: the prevailing practice is adapting off-the-shelf foundation models to downstream tasks through data-specific fine-tuning, rather than undertaking pre-training from scratch. Substantial random noise primarily emerges in industrial-scale datasets, but pre-training on such massive datasets exceeds our computational resources. Furthermore, our choice to fine-tune the downstream task head strictly follows [17], aiming to enhance model performance under the assumption of black-box model constraints. Additionally, it should be clarified that our core experimental validation resides in Table 1, while the additional experiments conducted on both LLMs and visual models serve as supplementary evidence to demonstrate the generalization capability of our proposed method - this constitutes an extended validation rather than the primary focus of this paper.
4.1 Method
In the preceding analysis, we demonstrate that the population-level loss function is only marginally affected by random noise. However, during the SGD training process, its presence introduces certain noise into the gradients. Prior studies [33, 34] have shown that artificially added gradient noise can hurt the model’s generalization. Therefore, we propose explicitly enhancing the denoising capabilities of the downstream task head by aligning local gradients.
@@ 4.1 Method
Specifically, let $C$ denote the number of classes in the downstream task, and let $g_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ represent the linear or MLP classification head parameterized by $\theta$. Let $t^{*}$ be the feature extracted by $h^{*}$, $t$ be the feature extracted by $h$, and $y$ be the corresponding label.
@@
Let $\ell(\hat{y},y)$ be the loss function(typically cross-entropy), and $\mathcal{L}_{ce}(\mathcal{D},g_{\theta})=\mathbb{E}_{(t,y)\sim\mathcal{D}}\ell(g_{\theta}(t),y)$ be the population-level loss where $\mathcal{D}$ represents the joint distribution of downstream features and labels. Due to the additional randomness introduced by $h$ as a result of noise, $t$ can be viewed as $t^{*}$ perturbed by minor disturbances. If both $t^{*}$ and $t$ were known, their distribution could be aligned to achieve denoising. However, in practical applications, it is challenging to obtain $t^{*}$. To construct contrastive sample pairs without $t^{*}$, we add Gaussian noise to $t$ to obtain $\hat{t}$:
Table 4: Evaluation of our method combined with SOTA fine-tuning techniques utilizing BERT-Large ( cited from [35]).
@@
Table 5: Ablation Study. To investigate the effects of reducing $\mathcal{L}_{gm}$, we separately reduce the norm or increase the cosine similarity.
@@
$\hat{t}=t+\gamma\cdot\delta$ (5)
4.2 Theoretical Analysis
Proposition 2.
Proposition 2 demonstrates that $\mathcal{L}_{gm}$ is closely associated with the smoothness of the loss function in both the parameter space and the input space. As a flat minima is widely acknowledged to benefit the generalization of neural networks [36, 37], it explains the effectiveness of $\mathcal{L}_{gm}$. The final loss function is:
$\mathcal{L}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{gm}\vspace{-2mm}$ (8)
4.3 Experiments
We first conduct extensive experiments using trained GPT-2 models. Then, to further validate the novelty and effectiveness of the LGM loss, we conduct additional experiments using Llama-3 [38] and vision models. These experiments are intended to showcase the generalizability of our approach beyond the specific context of GPT-2, demonstrating its applicability across different types of models and tasks. While these experiments enrich our study, they are not the core focus but rather supplementary evidence supporting the broader applicability of our proposed solution. Details can be found in Appendix E.
@@ 4.3 Experiments
0.001 0.05 76.48 78.42
0.05 0.1 76.81 79.21
0.1 0.15 76.59 79.37
We validate the performance of $\mathcal{L}_{gm}$ on models pre-trained with noisy data using four commonly used classification datasets: SST-2, SST-fine, 20newsgroup, and CR. The training hyperparameters follow those of [24], where $\gamma=0.01$ and $\lambda=0.15$ apply to all four experiments. In line with the approach described by [17], we freeze the model parameters and only fine-tune a linear or MLP classifier head. As shown in Table 1, our model achieves competitive results without reaching the number of training iterations of GPT-2, and $\mathcal{L}_{gm}$ consistently boosts performance.
(a) No regularization.
(b) With $L^{2}$ regularization.
(c) With $\mathcal{L}_{gm}$.
@@ 5 Conclusion
In this paper, we investigate the random noise present in language model pre-training datasets, which is inevitable in real-world scenarios but receives little attention. We pre-train multiple GPT-2 models under varying noise levels and find that random noise has a minor impact on the pre-training loss. We then provide a theoretical explanation for this phenomenon and discover that our theory can elucidate the success of multilingual and multimodal models. Interestingly, we observe that slight noise can sometimes enhance a model’s generalization ability. Then, building on the noisy model learning setup, we propose a novel local gradient matching loss. Extensive experiments across multiple datasets in both language and vision tasks, as well as with various backbone models, validate the effectiveness of our proposed method. We hope this work inspires more researchers to focus on data-centric AI.
@@ B.1 Explanation of Equation (1)
First, define a new probability measure $P_{3}$ as $P_{3}(A)=\alpha P_{1}(A)+(1-\alpha)P_{2}(A)$ for any measurable set $A\subseteq\mathcal{M}$. Here, $P_{3}$ is a convex combination of $P_{1}$ and $P_{2}$, and thus $P_{3}$ is also a valid probability measure [39].
@@ B.3 Justification of Proposition 1
(1) Numerous studies on compressing large language models, such as pruning [40, 41, 42], quantization [43, 44, 45], and distillation [46, 47], have demonstrated that there exists a significant amount of redundancy within the parameters of large models. Therefore, we could first train a model on $P^{c}$ and then compress it, fine-tuning the surplus parameters on $P^{n}$. This approach would allow us to improve $p_{n}$ without altering $p_{c}$.
(2) A small proportion of data corresponding to $P^{n}$ can be introduced into $P^{c}$, making sure that $\alpha$ is extremely small. According to domain adaptation theory [23], this would only slightly increase $\mathcal{L}_{ntp}(P^{c})$. However, existing results [48, 28, 29] indicate that pre-trained models like BERT or GPT on English text can exhibit strong multilingual capabilities with just a very limited amount of data. Consequently, compared to a model trained solely on $P^{c}$, the resulting model has a minor difference in $p_{c}$ but a relatively higher $p_{n}$.
@@
Definition 1 ($\beta$-smooth [49]).
@@
where $B(0,\rho)=\{\delta^{\prime}:||\delta^{\prime}||_{2}<\rho\}$ is an open ball.
@@ Appendix C Detailed Related Works
Data selection for language model training. LLMs [51, 52, 53, 54, 55, 56, 57] have fundamentally change the landsape of current AI research [58, 59, 60, 61, 62, 63]. Large text corpora form the backbone of language models, with data quality being fundamental to their success and safety [64, 65]. [2] conducted a systematic analysis of open-source text datasets such as The Pile [21] (used to train Pythia), C4 [66] (used to train T5) and RedPajama (used to train LLaMA), revealing that they contain a significant amount of duplicate, toxic, synthetic, and low-quality content. Therefore, it is of great importance to thoroughly understand the impact of low-quality data within these pre-training datasets on the model’s performance, reliability, and safety. [4, 67] systematically investigated the effect of low-quality data and found that such data can significantly reduce the model’s knowledge capacity, sometimes by up to 20 times. Another research direction primarily focuses on the synthetic data of large language models, specifically examining the impacts of using data generated by LLMs for recursive training. The study by [6] was the first to explore this issue and introduced the concept of "model collapse", indicating that recursive training can lead to the loss of information in tail tokens, ultimately resulting in the model producing nonsensical content. [7] mainly provided a theoretical explanation for why model collapse occurs, supporting their arguments with experimental evidence. Consequently, the importance of data selection cannot be overstated. Given that data selection is an NP-hard problem in terms of combinatorial optimization [68], numerous heuristic algorithms have been proposed to expedite the process. [1] provided a comprehensive study on pre-training data selection and optimal ratios, offering practical recommendations. [69] proposed dataset pruning, an approach that assesses the impact of omitting training samples on model generalization and creates a minimal training subset with a controlled generalization gap. [70] evaluated the impact of individual training samples on the dynamics of GPT model training. [71] introduced the Instruction-Following Difficulty metric to assess the quality of instruction-tuning data. [5] employed importance resampling for data selection. [72, 73] advocated for optimizing data composition and diversity. Despite these notable studies on data selection, they generally acknowledge that dataset noise degenerates model performance but lack a detailed understanding of how and to what extent, particularly in the case of random noise which is inevitable in large-scale datasets. Although [74] investigated the influence of gibberish input, the random noise within the pre-training dataset is still underexplored. This paper aims to bridge the gap.
@@
Learning from Noisy Distributions. The majority of machine learning algorithms assume that training and test samples are independently and identically distributed (i.i.d.), a condition that is often not met in real-world scenarios. For instance, LLMs are pre-trained on datasets with all kinds of noise while their performance is evaluated by the user whose distribution is usually clean and meets real-world scenarios, which violates the i.i.d. assumption. Domain adaptation [75, 76, 77] addresses this issue when the distribution of the training data differs from that of the test data. Although domain adaptation methods attempt to reduce the statistical distribution discrepancy [78, 79] or employ adversarial training [80, 81] to minimize the gap between source and target domains, they typically require access to unlabeled test data under a semi-supervised learning setup, which is impractical for LLM training. Another reason domain adaptation cannot be directly applied here is that domain adaptation theory [23] focuses on the performance of a model trained on one distribution when it is applied to another different but related distribution. This kind of bounds can be easily derived by Lemma 1. However, what we aim to investigate here is the extent of performance loss when comparing a model trained on one distribution (noisy dataset) to a model trained on another distribution (clean dataset).
@@
Apart from domain adaptation, there has been extensive research directly investigating noisy training sets. Noisy label learning [82, 83] have explored the impact of incorrect labels on model performance. Regarding input feature noise, [84] added perturbations to individual image inputs to enhance model interpretability, and [85] added white noise image into the training dataset to tackle the class imbalance problem. However, most of these efforts have concentrated on image classification and do not consider the pre-training paradigm.
@@
Fine-tuning Pre-trained Models. The approach of initially pre-training model weights on large-scale datasets and subsequently fine-tuning them with downstream data has become the de facto standard in the fields of computer vision [86, 15, 87, 88], natural language processing [25, 89, 90, 91] and audio processing [16, 92, 93]. For instance, [94, 35] proposed enhancing the performance of models by increasing their resistance to minor perturbations in intermediate layers. Meanwhile, [95] improved model robustness by adding regularization terms. Besides full-parameter fine-tuning, numerous parameter-efficient fine-tuning algorithms have been extensively studied. [96] introduced adapters into the original model architecture, optimizing only these parameters during fine-tuning. [97, 98] efficiently fine-tuned CLIP models [99] using learnable soft prompts. [100] optimized models through learning low-rank residual weights. These methods achieved performance close to that of full-parameter fine-tuning while maintaining the generalization ability of the original models. However, they all require access to the model’s weights and loading them into GPU memory, which can be challenging for today’s large models, especially when state-of-the-art models’ parameters are not publicly available. Therefore, in this paper, we follow the NML setup and explore efficient ways to fine-tune the downstream task head under a black-box scenario.
@@
Implicit Regularization and Sharpness-aware Minimization. Achieving good generalization in neural networks optimized using gradient descent algorithms has long been a research focus in deep learning theory. [101] explored the properties of stochastic gradient descent (SGD), finding that SGD implicitly constrains the gradient norm. Based on this observation, Sharpness-aware minimization (SAM) [102, 103, 104, 105] improves generalization by incorporating the gradient norm as a regularization term. Our method can be seen as drawing inspiration from SAM but differs in that our optimization objective is the model’s resilience to input noise rather than seeking flat minima in the parameter space.